# Tunable absorption spectrum splitting in a pulse-driven three-level system

### v1

@@ III Results
We start by assessing the realization of the steady state in our pulse-driven system. Fig. 2 shows $\rho_{33}(t)$, the population of the highest excited state (|3)|3$\rangle$)| 3 ⟩ ) as a function of time, when the $(|1\rangle\leftrightarrow|2\rangle)$ transition is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.3$. Here, the pulses are due to a driving field of Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. $\rho_{33}(t)$ is shown for $\omega=0$ (solid green line), $\pi/2\tau$ (dashed blue line), $3\pi/2\tau$ (dashed red line). The figure clearly shows that after an initial non-trivial transient, the system settles into a steady state where the population for any frequency does not show any non-trivial change as a function of time. The rest of the data that we present throughout this paper corresponds to this steady state. Here and in what follows, we choose $\Gamma_{3}=0.2$ and $\Omega_{2}=1.0$ but we note that the same lineshape is obtained for $\Omega_{2}=0.1$ with the only difference that the peaks in the absorption spectrum have higher magnitudes for larger $\Omega_{2}$.
@@
Fig. 3 shows, for detuning $\Delta_{1}=0$, the absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition, or the population of the highest excited state as a function of frequency, when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with different values of the period $\tau$. The pulse is due to a driving field with Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. The figure shows results for inter-pulse delays, $\tau=0.1$ (black solid line), $\tau=0.2$ (dashed blue line), $\tau=0.3$ (dashed green line), $\tau=0.4$ (dotted red line). The spectrum features two symmetric main peaks separated by a frequency $\pi/\tau$ with a clear similarity with the Autler-Townes splitting obtained for a continuous resonant drive [19, 25, 24]. Note that there are additional peaks centered around frequencies that are integer multiples of ±$\pm\pi/2\tau$ but with exponentially suppressed spectral weights. These additional peaks with smaller spectral weight are visible in the figure for large $\tau$ values.
Figure 4: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau$ the pulse is due to a driving field of Rabi frequency $\Omega_{1}=100$ (solid back line), $\Omega_{1}=50$ (solid red line), $\Omega_{1}=25$ (dashed blue line), that is applied for time $t_{\pi}=\pi/\Omega_{1}$. The pulse period is $\tau=0.2$ and the detuning is $\Delta_{1}=0$ (top) and $\Delta_{1}=3$ (bottom).
Figure 5: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.2$ (top) and $\tau=0.3$ (bottom). The $\pi$ pulses are achieved with Rabi frequency $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$ (solid black line), $\Delta_{1}=3$ (dashed blue line), $\Delta_{1}=5$ (dashed green line), $\Delta_{1}=8$ (dotted red line).
@@
Next, we examine the dependence of this $|2\rangle\leftrightarrow|3\rangle$ absorption spectrum on the Rabi frequency of the pulses. This is equivalent to examining how this spectrum varies with the duration of the $\pi$ pulses. Fig. 4 shows this absorption spectrum for a pulse period $\tau=0.2$ for $\Delta_{1}=0$ (top) and $\Delta_{1}=3$ (bottom) for pulse Rabi frequency $\Omega_{1}=100$ (solid black line), $\Omega_{1}=50$ (solid red line), $\Omega_{1}=25$ (dashed blue line). One readily observes a weak dependence on the Rabi frequency. The lineshape is preserved even for relatively broad pulses ($t_{\pi}\sim\tau/2$). For finite detuning, we observe an asymmetry in the absorption spectrum and with more spectral weight shifted to the satellite peaks. This feature is emphasized in Fig. 5 that shows the absorption spectrum for different detunings. Here, $\tau=0.2$ (top panel) and $\tau=0.3$ (bottom panel). The $\pi$ pulses are achieved by a driving field with Rabi frequency is $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$ (solid black line), $\Delta_{1}=3$ (dashed blue line), $\Delta_{1}=5$ (dashed green line), $\Delta_{1}=8$ (dotted red line). Here, we observe the shift in the location of the peaks and their relative amplitudes as a function of the detuning with more pronounced asymmetry found in the results for larger inter-pulse delay.
@@ IV Conclusion
The Autler-Townes effect and its parent phenomenon, electromagnetically induced transparency, have been actively investigated since their discoveries. Their applications have been demonstrated in recent advances as quantum memories in quantum information processing. The Autler-Townes effect is readily explained by the dressed state picture that is also responsible for the Mollow triplet of resonance fluorescence occuring when a two-level system is driven on resonance by a strong incident field. Previous work has shown that when a two-level system is instead driven by a periodic sequence of $\pi$ pulses, the emission spectrum obtained has similarities with the Mollow triplet. This spectrum features a central peak at the pulse carrier frequency and two satellite peaks at ±$\pm\pi/\tau$ where $\tau$ is the inter-pulse delay of the sequence. In the present paper, we have evaluated, the absorption spectrum of the three-level system in the ladder configuration when the bottom two levels are driven by a periodic sequence of $\pi$ pulses, while the transition between the middle and the highest level is probed by a weak field. For a pulse carrier frequency that is resonant with the transition frequency of the bottom two levels, this absorption spectrum displays similarities with the Autler-Townes doublet. It features two main peaks at frequencies separated by $\pi/\tau$. This spectrum is found to have little dependence on the pulse carrier frequency and shows little change even for rather broad pulses (pulse duration $\sim\tau/2$). For finite detuning, the spectrum has an asymmetry that increases with the detuning. These results demonstrate the capacity to modulate the absorption spectrum of a three-level system with experimentally achievable pulse protocols and provide an alternative pathway for realizing quantum memories in these systems.

### v2

@@ III Results
We start by assessing the realization of the steady state in our pulse-driven system. Fig. 2 shows $\rho_{33}(t)$, the population of the highest excited state ($|3\rangle)$ as a function of time, when the $(|1\rangle\leftrightarrow|2\rangle)$ transition is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.3$. Here, the pulses are due to a driving field of Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. $\rho_{33}(t)$ is shown for $\omega=0$ (solid green line), $\pi/2\tau$ (dashed blue line), $3\pi/2\tau$ (dashed red line). The figure clearly shows that after an initial non-trivial transient, the system settles into a steady state where the population for any frequency does not show any non-trivial change as a function of time. The rest of the data that we present throughout this paper corresponds to this steady state. Here and in what follows, we choose $\Gamma_{3}=0.2$ and $\Omega_{2}=1.0$ but we note that the same lineshape is obtained for $\Omega_{2}=0.1$ with the only difference that the peaks in the absorption spectrum have higher magnitudes for larger $\Omega_{2}$.
@@
Fig. 3 shows, for detuning $\Delta_{1}=0$, the absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition, or the population of the highest excited state as a function of frequency, when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with different values of the period $\tau$. The pulse is due to a driving field with Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. The figure shows results for inter-pulse delays, $\tau=0.1$ (black dashed line), $\tau=0.2$ (dashed blue line), $\tau=0.3$ (dashed green line), $\tau=0.4$ (dotted red line). The spectrum features two symmetric main peaks separated by a frequency $\pi/\tau$ with a clear similarity with the Autler-Townes splitting obtained for a continuous resonant drive [19, 28, 27]. The separation between the main peaks is shown in Fig. 4 that presents this separation between the main peaks as a function of $1/\tau$. The figure clearly shows that this separation in the pulse-driven system is equal to $\pi/\tau$. Note that there are additional peaks centered around frequencies that are integer multiples of $\pm\pi/2\tau$ but with exponentially suppressed spectral weights. These additional peaks with smaller spectral weight are visible in the figure for large $\tau$ values. The solid lines in Fig. 3 show the absorption of the continuously driven system with frequency $\pi/\tau$ and all other parameters identical to the pulse-driven case for $\tau=0.1$ (black line), $\tau=0.2$ (blue line), $\tau=0.3$ (red line), $\tau=0.4$ (green line). Note the identical peak locations. The peak heights are different because of the relative values of the probing field Rabi frequency with respect to the driving field.
We show the requirement of the $\pi$ pulses in Fig. 5 where the absorption is shown for a periodic sequence of pulses with each pulses accomplishing a $0.8\pi$ rotation (dashed green line), $0.9\pi$ rotation (dashed red line), $1.0\pi$ rotation (solid black line), $1.1\pi$ rotation (dashed blue line), $1.2\pi$ rotation (dashed purple line), with period $\tau=0.3$ and $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$. We see that as the pulses are tuned away from a $\pi$ rotation, the peak separations are adjusted away from the $\pi/\tau$ value.
@@
Next, we examine the dependence of this $|2\rangle\leftrightarrow|3\rangle$ absorption spectrum on the Rabi frequency of the pulses. This is equivalent to examining how this spectrum varies with the duration of the $\pi$ pulses. Fig. 6 shows this absorption spectrum for a pulse period $\tau=0.2$ for $\Delta_{1}=0$ (top) and $\Delta_{1}=3$ (bottom) for pulse Rabi frequency $\Omega_{1}=100$ (solid black line), $\Omega_{1}=50$ (solid red line), $\Omega_{1}=25$ (dashed blue line). One readily observes a weak dependence on the Rabi frequency. The lineshape is preserved even for relatively broad pulses ($t_{\pi}\sim\tau/2$). For finite detuning, we observe an asymmetry in the absorption spectrum and with more spectral weight shifted to the satellite peaks. This feature is emphasized in Fig. 7 that shows the absorption spectrum for different detunings. Here, $\tau=0.2$ (top panel) and $\tau=0.3$ (bottom panel). The $\pi$ pulses are achieved by a driving field with Rabi frequency $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$ (solid black line), $\Delta_{1}=3$ (dashed blue line), $\Delta_{1}=5$ (dashed green line), $\Delta_{1}=8$ (dotted red line). Here, we observe the shift in the location of the peaks and their relative amplitudes as a function of the detuning with more pronounced asymmetry found in the results for larger inter-pulse delay.
@@ IV Conclusion
The Autler-Townes effect and its parent phenomenon, electromagnetically induced transparency, have been actively investigated since their discoveries. Their applications have been demonstrated in recent advances as quantum memories in quantum information processing. The Autler-Townes effect is readily explained by the dressed state picture that is also responsible for the Mollow triplet of resonance fluorescence occuring when a two-level system is driven on resonance by a strong incident field. Previous work has shown that when a two-level system is instead driven by a periodic sequence of $\pi$ pulses, the emission spectrum obtained has similarities with the Mollow triplet. This spectrum features a central peak at the pulse carrier frequency and two satellite peaks at $\pm\pi/\tau$ where $\tau$ is the inter-pulse delay of the sequence. In the present paper, we have evaluated, the absorption spectrum of the three-level system in the ladder configuration when the bottom two levels are driven by a periodic sequence of $\pi$ pulses, while the transition between the middle and the highest level is probed by a weak field. For a pulse carrier frequency that is resonant with the transition frequency of the bottom two levels, this absorption spectrum displays similarities with the Autler-Townes doublet. It features two main peaks at frequencies separated by $\pi/\tau$. This spectrum is found to have little dependence on the pulse carrier frequency and shows little change even for rather broad pulses (pulse duration $\sim\tau/2$). For finite detuning, the spectrum has an asymmetry that increases with the detuning. These results demonstrate the capacity to modulate the absorption spectrum of a three-level system with experimentally achievable pulse protocols and provide an alternative pathway for realizing quantum memories in these systems.